# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 8. Generalised double-wheels
To deal with the generalised double-wheel outcome of Theorem 1, we first introduce a particular cycle-decomposition in Section 8.1 for analysing the wheel-structure obtained from two tetra-separations that cross with vertex-centre of size two. Section 8.2 introduces machinery that exploits the structure of this cycle-decomposition to find independent paths. Section 8.3 employs this machinery to show that all non-trivial bags of the cycle-decomposition give rise to totally-nested tetra-separations (8.27). Section 8.4 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1 and deduces that its torso is a generalised double-wheel (8.37). Finally, Section 8.5 characterises the special case where additionally all tetra-separations of the graph are crossed (8.41).
@@ 8.1. Block-bagel
(i). The bag $V_{t}$ is unique since $G$ is 4-connected. It follows from the degree-condition that $G[V_{t}]$ must be 2-connected.
(ii). The first part is immediate. For the ‘moreover’-part we use that $e$ is not dangling by the (Crossing Lemma). (4.1). ∎
@@
The bag $V_{t}$ and node $t$ in Lemma 8.2 are said to represent the $L$-link in the $K$-block-decomposition and in the $K$-block-path.
@@
    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if their bags $G[V_{t}]$ and $G[V_{t^{\prime}}]$ are 2-connected;
@@
    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node $t$ of $O$ with $t=o(L)$ for some link $L$, we let $G_{t}$ be the $K_{2}$-subgraph of $G$ from Lemma 8.2 (ii) that represents the $L$-link in the $K$-block-decompositions for the two corners $K$ adjacent to $L$. The pair $(O,(G_{t})_{t\in V(O)})$ is the block-bagel induced by $(A,B)$ and $(C,D)$. It really is a block-bagel of $G-Z^{\prime}$.
The following setting is supported by Figure 8.1.
@@
Let $(A,B)$ and $(C,D)$ be two tetra-separations of a 4-connected graph $G$ that cross with all links of size one. Let $Z^{\prime}$ be the centre, which consists of two vertices $z^{\prime}_{1},z^{\prime}_{2}$. Let $\mathcal{O}=(O,(G_{t})_{t\in V(O)})$ be the Block-bagel induced by $(A,B)$ and $(C,D)$. Write $O=t(0),t(1),\ldots,t(m),t(0)$ where indices are treated as elements of $\mathbb{Z}_{m+1}$. For every node $t\in V(O)$, we let $U_{t}^{\prime}:=V(G_{t})\cup Z^{\prime}$ and $W_{t}^{\prime}:=Z^{\prime}\cup\bigcup_{s\in V(O-t)}V(G_{s})$, and we let $(U_{t},W_{t})$ be the right-left-shift of $(U_{t}^{\prime},W_{t}^{\prime})$.
@@
Figure 8.1. The (Block-bagel setting). (8.4). The vertices $z^{\prime}_{1}$ and $z^{\prime}_{2}$ are in the centre. Left: the arrangement of the bags. Center: The 4-separation+ $(U^{\prime},W^{\prime})$. The vertex $x^{\prime}$ is red, the vertex $y^{\prime}$ is green. Right: The right-left-shift $(U,W)$ of $(U^{\prime},W^{\prime})$. Here $x$ is the red edge and $y$ is the green vertex. One of the $z_{i}$ is a vertex, the other is the blue edge.
@@ 8.2. N-Paths lemmata
Let $K$ be an arbitrary corner. If the $K$-block-decomposition has a 2-connected bag, then we are done. Otherwise, the $K$-shrimp is a path. The corner $K$ contains a vertex $v$ by Corollary 8.3. Then $v$ is the adhesion-vertex of a helpful edge that belongs to the corner $K$. ∎
@@
The following two lemmata are supported by Figure 8.2.
@@
Figure 8.2. The neighbours $w_{i}$ found by Lemma 8.8 are blue. The paths found by Lemma 8.9 are red.
@@
Assume the  (Block-bagel setting). (8.4). Let $P\subseteq O$ be a path of length $\geqslant 2$ with end-nodes $r$ and $s$. Let $u\in V(G_{r})$ and $v\in V(G_{s})$. Then there is a $u$$v$ path in $\bigcup_{t\in V(P)}G_{t}$ that internally avoids the adhesion-vertices of the edges not in $P$.
@@
Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$. We assume without loss of generality that $P=t(1)\,t(2)\ldots t(k)$. For each $i\in\{1,2,\ldots,k\}$, we construct a path $P_{i}$ in $G_{t(i)}$, as follows. For $i=1$ we claim that $P_{i}$ can be chosen to be a $u$$u_{1}$ path with interior in $G_{t(1)}-u_{0}$. This is clearly possible if the bag $G_{t(1)}$ is a $K_{2}$. Otherwise $G_{t(1)}$ is 2-connected. Then we find two independent $u$$u_{1}$ paths in $G_{t(1)}$, and at least one of them internally avoids $u_{0}$. Similarly, if $i=k$ then we let $P_{i}$ be a $u_{k-1}$$v$ path with interior in $G_{t(k)}-u_{k}$. Otherwise, if $2\leqslant i\leqslant k$, then we let $P_{i}$ be a $u_{i-1}$$u_{i}$ path in $G_{t(i)}$. Then the concatenation $P:=P_{1}P_{2}\ldots P_{k}$ is a desired $u$$v$ path. ∎
@@
Assume the  (Block-bagel setting). (8.4). Let $u$ and $v$ be the two adhesion-vertices in a bag $G_{t}$. Assume that either $t$ is helpful or $u$ is the adhesion-vertex of a helpful edge. Then there exist two independent paths $P_{1},P_{2}$ with all internal vertices in $\mathring{G}_{t}$ such that $P_{1}$ is a $u$$z_{1}^{\prime}$ path and $P_{2}$ is a $u$$v$ path.
@@
If $u$ is the adhesion-vertex of a helpful edge, then $uz_{1}^{\prime}\in E(G)$ by Lemma 8.7, and $uv\in E(G)$ since $G_{t}$ is a $K_{2}$. Then we let $P_{1}:=uz_{1}^{\prime}$ and $P_{2}:=uv$. Otherwise, $t$ is helpful. By Lemma 8.8, $z^{\prime}_{1}$ has a neighbour $w$ in $\mathring{G}_{t}$. Since $G_{t}$ is 2-connected, it contains two independent paths $Q_{1},Q_{2}$ such that $Q_{1}$ is a $u$$w$ path and $Q_{2}$ is a $u$$v$ path. Then we obtain $P_{1}$ from $Q_{1}$ by appending $z^{\prime}_{1}$, and we let $P_{2}:=Q_{2}$. ∎
The following lemma is supported by Figure 8.3.
@@
Assume the  (Block-bagel setting). (8.4). Let $P$ be an $r$$s$ path in $O$. Assume that either the first edge $e$ or the second node of $P$ is helpful. Assume that $P$ contains at least two helpful elements distinct from $r$. Let $u$ be the adhesion-vertex of the edge $e$. Then there exist two independent $u$$z_{1}^{\prime}$ paths in $G$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$.
@@
Without loss of generality, $P=t(0)\,t(1)\ldots t(k)$ with $r=t(0)$ and $s=t(k)$. By assumption, either the edge $e=t(0)\,t(1)$ or the second node $t(1)$ is helpful. Let $u_{i}$ denote the adhesion-vertex of $t(i)\,t(i+1)$. In particular, $u=u_{0}$. By the (Two-Fan Lemma). (8.11), there are two independent paths $P_{1},Q_{1}$ with interior in $\mathring{G}_{t(1)}$ such that $P_{1}$ is a $u_{0}$$z_{1}^{\prime}$ path and $Q_{1}$ is a $u_{0}$$u_{1}$ path.
@@ 8.3. Totally-nested tetra-separations from good bags
The following setting is supported by Figure 8.4.
@@
Assume the (Block-bagel setting). (8.4). Let $U^{\prime}:=U_{t(0)}^{\prime}$ and $W^{\prime}:=W_{t(0)}^{\prime}$. Let $x^{\prime}$ be the adhesion-vertex of $t(-1)\,t(0)$ and let $y^{\prime}$ be the adhesion-vertex of $t(0)\,t(1)$. Observe that $S(U^{\prime},W^{\prime})=\{x^{\prime},y^{\prime},z_{1}^{\prime},z_{2}^{\prime}\}$. Let $(U,W)$ be the right-left-shift of $(U^{\prime},W^{\prime})$. Let $x,y,z_{1},z_{2}\in S(U,W)$ correspond to $x^{\prime},y^{\prime},z^{\prime}_{1},z^{\prime}_{2}$, respectively. [^3] So $x$ is incident or equal to $x^{\prime}$, and so on. Let $x^{W}$ denote the unique vertex in $\hat{x}\cap W$, and define $y^{W}$ and $z_{i}^{W}$ similarly.
@@
We combine the (Helpful Lemma). (8.6) with Lemma 8.7 and Lemma 8.8, and we use 5.9 for $W^{\prime}\smallsetminus U^{\prime}\subseteq W\smallsetminus U$. ∎
@@
(i) $\Rightarrow$ (ii). Suppose first that $G_{t(0)}$ is good. By Corollary 8.3, each of the four corners contains a vertex. Since $G_{t(0)}$ is 2-connected, it is included in $A\cap C$, say, so every corner except $AC$ has its vertex contained in $W^{\prime}\smallsetminus U^{\prime}$. Hence $|W^{\prime}\smallsetminus U^{\prime}|\geqslant 3$. The 4-separation $(U^{\prime},W^{\prime})$ trivially satisfies the matching-condition. If the bag $G_{t(0)}$ is not a triangle, then it has $\geqslant 4$ vertices, so $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. In this case, $(U,W)$ is a tetra-separation by 5.10. Otherwise, $G_{t(0)}$ is a triangle. Then some neighbouring bag, say $G_{t(1)}$, is a $K_{2}$. Hence $|U^{\prime}\smallsetminus W^{\prime}|=1$ and the vertex $y^{\prime}$ has exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$. Again by 5.10, $(U,W)$ is a tetra-separation.
Suppose now that all three bags $G_{t(-1)},G_{t(0)},G_{t(1)}$ are $K_{2}$’s. Then $G_{t(0)}$ meets at most two corners, so the vertices in the other two corners lie in $W^{\prime}\smallsetminus U^{\prime}$, giving $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. Both $x^{\prime}$ and $y^{\prime}$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$ since both $G_{t(-1)}$ and $G_{t(1)}$ are $K_{2}$’s. Hence $(U,W)$ is a tetra-separation by 5.10.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). As (i) fails, $G_{t(0)}$ either is a $K_{2}$ neighboured by at least one 2-connected bag, or $G_{t(0)}$ is a triangle neighboured by 2-connected bags. Suppose first that $G_{t(0)}$ is a $K_{2}$ neighboured by a 2-connected bag, say $G_{t(1)}$. Then $y^{\prime}$ has two neighbours in $G_{t(1)}$, which are also contained in $W^{\prime}\smallsetminus U^{\prime}$. The two vertices $z^{\prime}_{i}$ in the centre have $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16. So only $x^{\prime}$ might have $\leqslant 1$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=0$, so $(U,W)$ is no tetra-separation by 5.10.
Suppose now that $G_{t(0)}$ is a triangle with both neighbouring bags 2-connected. Then, as above, all vertices in $S(U^{\prime},W^{\prime})$ have $\geqslant 2$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=1$, so $(U,W)$ is no tetra-separation by 5.10. ∎
@@
Assume the  (Block-bagel bag setting). (8.14). Let $t(i)\,t(i+1)$ be the edge of $O$ that has $y^{W}$ as adhesion-vertex. Then either the edge $t(i)\,t(i+1)$ or the node $t(i+1)$ is helpful.∎
@@
For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows. If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma 8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure 8.5. If $h$ is a helpful node, then we use the (Bag-Path Lemma). (8.9) to let $P_{h}$ by a $z^{\prime}_{1}$$z^{\prime}_{2}$ path avoiding the edge $z^{\prime}_{1}z^{\prime}_{2}$ (if it exists) and with interior in $\mathring{G}_{h}$.
All paths $P_{h}$ are independent. By the (Helpful Lemma). (8.6), there are at least four of them. At most one of them meets the bag $G_{t(0)}$. By Lemma 8.16, both vertices $z^{\prime}_{i}$ in the centre are contained in $W$. Hence at least three of the paths $P_{h}$ are independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}-z^{\prime}_{1}z^{\prime}_{2}$. ∎
Figure 8.5. The red paths are constructed in Lemma 8.19.
The following lemma is supported by Figure 8.6.
@@
(6) $\geqslant 3$ elements of $O-t(0)$ are helpful.
Let $f=t(j)\,t(j+1)$ be the edge on $O$ that has $y^{W}$ as adhesion-vertex. Then $f$ or $t(j+1)$ is helpful by 8.18. If $t(j)$ is helpful, then $j=0$, and $j=1$ otherwise. Let $P$ be the path in $O$ that starts with $t(j),t(j+1)$ and has two helpful elements distinct from $t(j)$ and is minimal with these properties. The path $P$ is well-defined since $O-t(0)$ has $\geqslant 3$ helpful elements by 6. Moreover, the end $s$ of $P$ is distinct from $t(0)$. By the (Two-Paths Lemma). (8.12), we find two independent $y^{W}$$z_{1}^{\prime}$ paths $Q_{1},Q_{2}$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$. By construction, the paths $Q_{1}$ and $Q_{2}$ are contained in $G[W]$.
Now let $e=t(i)\,t(i-1)$ be the edge on $O$ that has $x^{W}$ as adhesion-vertex. Then $e$ or $t(i-1)$ is helpful by 8.18. And if $t(i)$ is helpful, then $i=0$. By the (One-Path Lemma). (8.10), we find an $x^{W}$$z_{1}^{\prime}$ path $Q_{3}$ with interior in $\mathring{G}_{t(i-1)}$. Once more, $Q_{3}\subseteq G[W]$ by construction.
It remains to show that the three paths $Q_{i}$ are independent. We already know that $Q_{1}$ and $Q_{2}$ are independent, so it remains to show that $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. If the edge $e$ is helpful, then it is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, $e$ does not lie on $P$ by 6. Since the bag $G_{t(i-1)}$ is a $K_{2}$, it has empty interior. So $Q_{3}=x^{W}z^{\prime}_{1}$ avoids $\mathring{G}_{s}$, even if $s=t(i-1)$, so $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. Otherwise $t(i-1)$ is helpful. Then $t(i-1)$ is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, both edges on $O$ incident to $t(i-1)$ are unhelpful. So $t(i-1)$ does not lie on $P$ by 6. Hence $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. ∎
Figure 8.6. The paths constructed in Lemma 8.20. $P_{1}$ is yellow, $P_{2}$ is blue, $P_{3}$ is green.
@@
Let $G$ be a graph, let $u\in U\subseteq V(G)$ and $v\in V\subseteq V(G)$. Every $u$$v$ path in $G$ contains a $U$$V$ path.∎
@@
We show that $G[W\smallsetminus U]$ is connected. Without loss of generality, assume that the interior of $G_{t(0)}$ is contained in the corner $AC$. Let $v$ be a vertex in the corner $BD$, which exists by Corollary 8.3. Let $t(i)$ be a node of $O$ with $v\in G_{t(i)}$.
We claim that every vertex $w$ in $W$ is connected to $v$ in $G[W\smallsetminus U]$. Since $G_{t(i)}\subseteq G[W\smallsetminus U]$, every vertex in $G_{t(i)}$ is connected to $v$ in $G[W\smallsetminus U]$. If $w$ does not lie in the centre $Z^{\prime}$, then we are done by the (One-Path Lemma). (8.10). Otherwise $w=z^{\prime}_{1}$, say. Then $w$ has a neighbour in $W\smallsetminus Z^{\prime}$ by Lemma 8.19, which we already know is linked to $v$ in $G[W\smallsetminus U]$. Hence $(U,W)$ is half-connected. ∎

### v2

@@ 8. Generalised double-wheels
To deal with the generalised double-wheel outcome of Theorem˜1, we first introduce a particular cycle-decomposition in Section˜8.1 for analysing the wheel-structure obtained from two tetra-separations that cross with vertex-centre of size two. Section˜8.2 introduces machinery that exploits the structure of this cycle-decomposition to find independent paths. Section˜8.3 employs this machinery to show that all non-trivial bags of the cycle-decomposition give rise to totally-nested tetra-separations (˜8.27). Section˜8.4 shows that these totally-nested tetra-separations form a splitting star as in Theorem˜1 and deduces that its torso is a generalised double-wheel (˜8.37). Finally, Section˜8.5 characterises the special case where additionally all tetra-separations of the graph are crossed (˜8.41).
@@ 8.1. Block-bagel
˜(i). The bag $V_{t}$ is unique since $G$ is 4-connected. It follows from the degree-condition that $G[V_{t}]$ must be 2-connected.
˜(ii). The first part is immediate. For the ‘moreover’-part we use that $e$ is not dangling by the (Crossing Lemma). (4.1). ∎
@@
The bag $V_{t}$ and node $t$ in Lemma˜8.2 are said to represent the $L$-link in the $K$-block-decomposition and in the $K$-block-path.
@@
[…]
@@
[…]
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node $t$ of $O$ with $t=o(L)$ for some link $L$, we let $G_{t}$ be the $K_{2}$-subgraph of $G$ from Lemma˜8.2 ˜(ii) that represents the $L$-link in the $K$-block-decompositions for the two corners $K$ adjacent to $L$. The pair $(O,(G_{t})_{t\in V(O)})$ is the block-bagel induced by $(A,B)$ and $(C,D)$. It really is a block-bagel of $G-Z^{\prime}$.
The following setting is supported by Figure˜8.1.
@@
Let $(A,B)$ and $(C,D)$ be two tetra-separations of a 4-connected graph $G$ that cross with all links of size one. Let $Z^{\prime}$ be the centre, which consists of two vertices $z^{\prime}_{1},z^{\prime}_{2}$. Let $\mathcal{O}=(O,(G_{t})_{t\in V(O)})$ be the Block-bagel induced by $(A,B)$ and $(C,D)$. Write $O=t(0),t(1),\ldots,t(m),t(0)$ where indices are treated as elements of $\mathbb{Z}_{m+1}$. For every node $t\in V(O)$, we let $U_{t}^{\prime}:=V(G_{t})\cup Z^{\prime}$ and $W_{t}^{\prime}:=Z^{\prime}\cup\bigcup_{s\in V(O-t)}V(G_{s})$, and we let $(U_{t},W_{t})$ be the right-left-reduction of $(U_{t}^{\prime},W_{t}^{\prime})$.
@@
Figure 8.1. The (Block-bagel setting). (8.4). The vertices $z^{\prime}_{1}$ and $z^{\prime}_{2}$ are in the centre. Left: the arrangement of the bags. Center: The 4-separation+ $(U^{\prime},W^{\prime})$. The vertex $x^{\prime}$ is red, the vertex $y^{\prime}$ is green. Right: The right-left-reduction $(U,W)$ of $(U^{\prime},W^{\prime})$. Here $x$ is the red edge and $y$ is the green vertex. One of the $z_{i}$ is a vertex, the other is the blue edge.
@@ 8.2. N-Paths lemmata
Let $K$ be an arbitrary corner. If the $K$-block-decomposition has a 2-connected bag, then we are done. Otherwise, the $K$-shrimp is a path. The corner $K$ contains a vertex $v$ by Corollary˜8.3. Then $v$ is the adhesion-vertex of a helpful edge that belongs to the corner $K$. ∎
@@
The following two lemmata are supported by Figure˜8.2.
@@
Figure 8.2. The neighbours $w_{i}$ found by Lemma˜8.8 are blue. The paths found by Lemma˜8.9 are red.
@@
Assume the  (Block-bagel setting). (8.4). Let $P\subseteq O$ be a path of length $\geqslant 2$ with end-nodes $r$ and $s$. Let $u\in V(G_{r})$ and $v\in V(G_{s})$. Then there is a uuvv path in $\bigcup_{t\in V(P)}G_{t}$ that internally avoids the adhesion-vertices of the edges not in $P$.
@@
[…]
@@
[…]
@@
If $u$ is the adhesion-vertex of a helpful edge, then $uz_{1}^{\prime}\in E(G)$ by Lemma˜8.7, and $uv\in E(G)$ since $G_{t}$ is a $K_{2}$. Then we let $P_{1}:=uz_{1}^{\prime}$ and $P_{2}:=uv$. Otherwise, $t$ is helpful. By Lemma˜8.8, $z^{\prime}_{1}$ has a neighbour $w$ in $\mathring{G}_{t}$. Since $G_{t}$ is 2-connected, it contains two independent paths $Q_{1},Q_{2}$ such that $Q_{1}$ is a uuww path and $Q_{2}$ is a uuvv path. Then we obtain $P_{1}$ from $Q_{1}$ by appending $z^{\prime}_{1}$, and we let $P_{2}:=Q_{2}$. ∎
The following lemma is supported by Figure˜8.3.
@@
[…]
@@
Without loss of generality, $P=t(0)\,t(1)\ldots t(k)$ with $r=t(0)$ and $s=t(k)$. By assumption, either the edge e=t(0)t(1)e=t(0)\,t(1) or the second node $t(1)$ is helpful. Let $u_{i}$ denote the adhesion-vertex of t(i)t(i+1)t(i)\,t(i+1). In particular, $u=u_{0}$. By the (Two-Fan Lemma). (8.11), there are two independent paths $P_{1},Q_{1}$ with interior in $\mathring{G}_{t(1)}$ such that $P_{1}$ is a $u_{0}$$z_{1}^{\prime}$ path and $Q_{1}$ is a $u_{0}$$u_{1}$ path.
@@ 8.3. Totally-nested tetra-separations from good bags
The following setting is supported by Figure˜8.4.
@@
Assume the (Block-bagel setting). (8.4). Let $U^{\prime}:=U_{t(0)}^{\prime}$ and $W^{\prime}:=W_{t(0)}^{\prime}$. Let $x^{\prime}$ be the adhesion-vertex of t(1)t(0)t(-1)\,t(0) and let $y^{\prime}$ be the adhesion-vertex of t(0)t(1)t(0)\,t(1). Observe that $S(U^{\prime},W^{\prime})=\{x^{\prime},y^{\prime},z_{1}^{\prime},z_{2}^{\prime}\}$. Let $(U,W)$ be the right-left-reduction of $(U^{\prime},W^{\prime})$. Let $x,y,z_{1},z_{2}\in S(U,W)$ correspond to $x^{\prime},y^{\prime},z^{\prime}_{1},z^{\prime}_{2}$, respectively. [^4] So $x$ is incident or equal to $x^{\prime}$, and so on. Let $x^{W}$ denote the unique vertex in $\hat{x}\cap W$, and define $y^{W}$ and $z_{i}^{W}$ similarly.
@@
We combine the (Helpful Lemma). (8.6) with Lemma˜8.7 and Lemma˜8.8, and we use ˜5.9 for $W^{\prime}\smallsetminus U^{\prime}\subseteq W\smallsetminus U$. ∎
@@
˜(i) $\Rightarrow$ ˜(ii). Suppose first that $G_{t(0)}$ is good. By Corollary˜8.3, each of the four corners contains a vertex. Since $G_{t(0)}$ is 2-connected, it is included in $A\cap C$, say, so every corner except $AC$ has its vertex contained in $W^{\prime}\smallsetminus U^{\prime}$. Hence $|W^{\prime}\smallsetminus U^{\prime}|\geqslant 3$. The 4-separation $(U^{\prime},W^{\prime})$ trivially satisfies the matching-condition. If the bag $G_{t(0)}$ is not a triangle, then it has $\geqslant 4$ vertices, so $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. In this case, $(U,W)$ is a tetra-separation by ˜5.10. Otherwise, $G_{t(0)}$ is a triangle. Then some neighbouring bag, say $G_{t(1)}$, is a $K_{2}$. Hence $|U^{\prime}\smallsetminus W^{\prime}|=1$ and the vertex $y^{\prime}$ has exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$. Again by ˜5.10, $(U,W)$ is a tetra-separation.
Suppose now that all three bags $G_{t(-1)},G_{t(0)},G_{t(1)}$ are $K_{2}$’s. Then $G_{t(0)}$ meets at most two corners, so the vertices in the other two corners lie in $W^{\prime}\smallsetminus U^{\prime}$, giving $|U^{\prime}\smallsetminus W^{\prime}|\geqslant 2$. Both $x^{\prime}$ and $y^{\prime}$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$ since both $G_{t(-1)}$ and $G_{t(1)}$ are $K_{2}$’s. Hence $(U,W)$ is a tetra-separation by ˜5.10.
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). As ˜(i) fails, $G_{t(0)}$ either is a $K_{2}$ neighboured by at least one 2-connected bag, or $G_{t(0)}$ is a triangle neighboured by 2-connected bags. Suppose first that $G_{t(0)}$ is a $K_{2}$ neighboured by a 2-connected bag, say $G_{t(1)}$. Then $y^{\prime}$ has two neighbours in $G_{t(1)}$, which are also contained in $W^{\prime}\smallsetminus U^{\prime}$. The two vertices $z^{\prime}_{i}$ in the centre have $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16. So only $x^{\prime}$ might have $\leqslant 1$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=0$, so $(U,W)$ is no tetra-separation by ˜5.10.
Suppose now that $G_{t(0)}$ is a triangle with both neighbouring bags 2-connected. Then, as above, all vertices in $S(U^{\prime},W^{\prime})$ have $\geqslant 2$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$. But $|U^{\prime}\smallsetminus W^{\prime}|=1$, so $(U,W)$ is no tetra-separation by ˜5.10. ∎
@@
Assume the  (Block-bagel bag setting). (8.14). Let t(i)t(i+1)t(i)\,t(i+1) be the edge of $O$ that has $y^{W}$ as adhesion-vertex. Then either the edge t(i)t(i+1)t(i)\,t(i+1) or the node $t(i+1)$ is helpful.∎
@@
For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows. If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure˜8.5. If $h$ is a helpful node, then we use the (Bag-Path Lemma). (8.9) to let $P_{h}$ by a $z^{\prime}_{1}$$z^{\prime}_{2}$ path avoiding the edge $z^{\prime}_{1}z^{\prime}_{2}$ (if it exists) and with interior in $\mathring{G}_{h}$.
All paths $P_{h}$ are independent. By the (Helpful Lemma). (8.6), there are at least four of them. At most one of them meets the bag $G_{t(0)}$. By Lemma˜8.16, both vertices $z^{\prime}_{i}$ in the centre are contained in $W$. Hence at least three of the paths $P_{h}$ are independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}-z^{\prime}_{1}z^{\prime}_{2}$. ∎
Figure 8.5. The red paths are constructed in Lemma˜8.19.
The following lemma is supported by Figure˜8.6.
@@
(6) $\geqslant 3$ elements of Ot(0)O-t(0) are helpful.
Let f=t(j)t(j+1)f=t(j)\,t(j+1) be the edge on $O$ that has $y^{W}$ as adhesion-vertex. Then $f$ or $t(j+1)$ is helpful by ˜8.18. If $t(j)$ is helpful, then $j=0$, and $j=1$ otherwise. Let $P$ be the path in $O$ that starts with $t(j),t(j+1)$ and has two helpful elements distinct from $t(j)$ and is minimal with these properties. The path $P$ is well-defined since Ot(0)O-t(0) has $\geqslant 3$ helpful elements by ˜6. Moreover, the end $s$ of $P$ is distinct from $t(0)$. By the (Two-Paths Lemma). (8.12), we find two independent $y^{W}$$z_{1}^{\prime}$ paths $Q_{1},Q_{2}$ with all interior vertices in $\bigcup_{t\in V(\mathring{P})}G_{t}\cup\mathring{G}_{s}$. By construction, the paths $Q_{1}$ and $Q_{2}$ are contained in $G[W]$.
Now let e=t(i)t(i1)e=t(i)\,t(i-1) be the edge on $O$ that has $x^{W}$ as adhesion-vertex. Then $e$ or t(i1)t(i-1) is helpful by ˜8.18. And if $t(i)$ is helpful, then $i=0$. By the (One-Path Lemma). (8.10), we find an $x^{W}$$z_{1}^{\prime}$ path $Q_{3}$ with interior in $\mathring{G}_{t(i-1)}$. Once more, $Q_{3}\subseteq G[W]$ by construction.
It remains to show that the three paths $Q_{i}$ are independent. We already know that $Q_{1}$ and $Q_{2}$ are independent, so it remains to show that $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. If the edge $e$ is helpful, then it is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, $e$ does not lie on $P$ by ˜6. Since the bag $G_{t(i-1)}$ is a $K_{2}$, it has empty interior. So $Q_{3}=x^{W}z^{\prime}_{1}$ avoids $\mathring{G}_{s}$, even if s=t(i1)s=t(i-1), so $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. Otherwise t(i1)t(i-1) is helpful. Then t(i1)t(i-1) is the first helpful element on $O$ following $t(0)$ in the direction of $e$. In particular, both edges on $O$ incident to t(i1)t(i-1) are unhelpful. So t(i1)t(i-1) does not lie on $P$ by ˜6. Hence $Q_{3}$ is independent from $Q_{1}$ and $Q_{2}$. ∎
Figure 8.6. The paths constructed in Lemma˜8.20. $P_{1}$ is yellow, $P_{2}$ is blue, $P_{3}$ is green.
@@
Let $G$ be a graph, let $u\in U\subseteq V(G)$ and $v\in V\subseteq V(G)$. Every uuvv path in $G$ contains a UUVV path.∎
@@
We show that $G[W\smallsetminus U]$ is connected. Without loss of generality, assume that the interior of $G_{t(0)}$ is contained in the corner $AC$. Let $v$ be a vertex in the corner $BD$, which exists by Corollary˜8.3. Let $t(i)$ be a node of $O$ with $v\in G_{t(i)}$.
We claim that every vertex $w$ in $W$ is connected to $v$ in $G[W\smallsetminus U]$. Since $G_{t(i)}\subseteq G[W\smallsetminus U]$, every vertex in $G_{t(i)}$ is connected to $v$ in $G[W\smallsetminus U]$. If $w$ does not lie in the centre $Z^{\prime}$, then we are done by the (One-Path Lemma). (8.10). Otherwise $w=z^{\prime}_{1}$, say. Then $w$ has a neighbour in $W\smallsetminus Z^{\prime}$ by Lemma˜8.19, which we already know is linked to $v$ in $G[W\smallsetminus U]$. Hence $(U,W)$ is half-connected. ∎